I have another test coming up and i know this is going to be on the test so i was trying to figure out how to do it but for the life of me i cant. so I was wondering if someone could explain to me specifically how to find how many 1/8 are in 3 and 1/2?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

I have another test coming up and i know this is going to be on the test so i was trying to figure out how to do it but for the life of me i cant. so I was wondering if someone could explain to me specifically how to find how many 1/8 are in 3 and 1/2?
-----------
3 1/2 = 7/2
= 28/8
--> 28

Answer by stanbon(75887)

how to find how many 1/8 are in 3 and 1/2?
-----------------
Convert 3 1/2 to an improper fraction:
3 1//2 = (2*3+1)/2 = 7/2
-------
Now divide 7/2 by 1/8
------------------------------
(7/2)/(1/8)
---
The math of the next step is to multiply
numerator and denominator by 8/1 so the
denominator becomes 1.
---
(7/2)*(8/1)
---
Multiply to get the final answer to your question:
56/2 = 28
---
Cheers,
Stan H.